OpenStudy (anonymous):

Solve the equation 3 ^5√(x+2)^3 +3=27

OpenStudy (anonymous):

sorry I cant draw it the 5 is like a cube root

OpenStudy (amorfide):

is that 3 to say question 3 or is that 3^5 or is that 3 multiplied by the rest?

OpenStudy (anonymous):

3 x the rest

OpenStudy (amorfide):

this?

OpenStudy (anonymous):

perfect yes

OpenStudy (amorfide):

first you want to move the +3 over, by subtracting 3 then you want to divide by 3, what do you get?

OpenStudy (anonymous):

then we divide both sides by 3

OpenStudy (amorfide):

then you want to make everything to the power of five on both sides, to get rid of the fifth root, then you want to cube root both sides to get rid of the cube, then solve for x

OpenStudy (anonymous):

so what would we get?

OpenStudy (amorfide):

subtract 3 gives us 24 divide by 3 gives us 8 raise everything to the 5th power gives us (x+2)^3=8^5 cuberoot both sides gives us x+2=2^5 x=32-2 x=30

OpenStudy (anonymous):

thank you